OpenStudy (anonymous):

Show that \[\sqrt{x+2}\rightarrow2 \] as \[x \rightarrow 2\]

OpenStudy (anonymous):

Are you looking for a hardcore proof?

OpenStudy (anonymous):

Yeah, from the definition of a limit. Should have absolute values etc.

OpenStudy (anonymous):

I could do it if it was, say a factorisable quadratic, but with the sqrt, i'm a bit confused.

OpenStudy (anonymous):

where x=0, sqrt of 2 is like 1.41 , where x = 1 sqrt of 3 is between 1.41 nd 2, just before x=2, the sqaure gets closer to 4 and the sqrt gets closer and closer to 2

OpenStudy (anonymous):

Alright, give me a sec

OpenStudy (anonymous):

but yeah thats ignoring negative values obviously

OpenStudy (anonymous):

Maybe even the first couple of steps re working with the sqrt function. I'd kind of like to figure most of it out myself.

OpenStudy (anonymous):

Let \[\epsilon > 0\] be given. We seek some \[ \delta > 0 \] such that \[ |x - 2| < \delta \text{ implies that } |\sqrt{x+2}-2| < \epsilon \] Let's restrict delta as follows: \[ \delta < 1 \] which means that \[ |x -2 | < 1 \rightarrow 1 < x < 3 \] Further, note that \[|x - 2| = |x+2 -4| = |\sqrt{x+2} -2 | \cdot |\sqrt{x+2} + 2| \] since \[ 1 < x < 3, |\sqrt{x+2} + 2| > \sqrt{3}+2 \] This means that \[ |x-2| > |\sqrt{x+2}-2| \cdot (\sqrt{3}+2) \] Therefore, let \[ \delta = min \left( 1 , \epsilon\cdot (\sqrt{3}+2) \right) \] So that if \[ |x-2| < \delta\] we have \[|x-2| = |\sqrt{x+2} -2| \cdot |\sqrt{x+2} -2| > |\sqrt{x+2}-2|\cdot (\sqrt{3}+2)\] \[ \text{ is less than } \epsilon \cdot (\sqrt{3}+2) \text{ which implies that } \] \[ |\sqrt{x+2} -2| < \epsilon \] whew. Hopefully I didn't make a typo...

OpenStudy (anonymous):

Almost definitely did though ;)

OpenStudy (anonymous):

Also oops, I didn't read your above message.... restrict yourself to the first few lines if you can haha

OpenStudy (anonymous):

Where did the x-2 come from?

OpenStudy (anonymous):

It's the limit as x approaches 2

OpenStudy (anonymous):

so |x-2| gets tiny

OpenStudy (anonymous):

So x gets within delta of 2, ok, gotcha.

OpenStudy (anonymous):

I get delta = min{1, (epsilon/(sqrt3+2))

OpenStudy (anonymous):

Yes you're right, I'm sorry. Told ya I probably made a typo somewhere....

OpenStudy (anonymous):

Actually, does it matter that you used the lower bound of x. If you use the 3, does it still hold?

OpenStudy (anonymous):

The way I wrote it, no, because of the > and < signs. Choosing delta < 1 was arbitrary though. Just for convenience.

OpenStudy (anonymous):

since \[ |x-2| = |\sqrt{x+2}-2| \cdot |\sqrt{x+2}+2| > |\sqrt{x+2}-2| \cdot (\sqrt{3}+2)\] we have that \[ |\sqrt{x+2}-2| < \frac{|x-2|}{\sqrt{3}+2} \] so if |x-2| < delta, where \[ \delta = \epsilon \cdot (\sqrt{3}+2) \] then \[| \sqrt{x+2}-2|< \frac{|x-2|}{\sqrt{3}+2} < \frac{\epsilon \cdot (\sqrt{3}+2)}{\sqrt{3}+2} = \epsilon \]